# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 4.2.2 Neural Machine Translation
Neural machine translation (NMT) refers to the automated process of translating text from one language to another language. Numerous deep learning models have been proposed for NMT which can be categorized into RNN-based and CNN-based models. One of the first successful RNN-based models is the encoder-decoder [44, 151]. The model consists of two connected subnetworks (the encoder and the decoder) for modelling the translation process as shown in Figure 24. The encoder reads the source sentence word by word and produces a fixed-length context vector (final hidden state). This process is known as source sentence encoding as shown in the figure. Given the context vector, the decoder generates the target sentence (translation) word by word. This modelling of the translation can be seen as a mapping between the source sentence to the target sentence via the intermediate context vector in the semantic space. The context vector represents the summary of the input sequence’s semantic meaning, providing a compressed representation that captures the essence of the source sentence. However, the compression process can sometimes result in the loss of information especially those early in the sequence. Bi-directional RNN may mitigate the loss of information by modelling the sequence in reverse order. However, the problem can still persist, particularly in cases where the input is a long sequence.
Figure 24: The architecture of an encoder-decoder [152].
@@
Attention mechanism was introduced to solve the problem of learning long input sequences [153]. Attention alleviates this issue by attending on different words of the input sequences when predicting the target sequences at each time step. Unlike the standard encoder-decoder model, attention derives the context vector from the hidden states of both the encoder and decoder, and the alignment between the source and target. This mechanism allows the model to focus on the important words, increasing the overall accuracy of the translation. Several alignment score functions have been proposed for calculating the attention weights. Some of the popular functions are additive [153], dot-product, location-based [154], and scaled dot-product [16]. The attention weights are calculated by attending to the entire hidden states of the encoder. This attention, also known as global attention, is computationally expensive. Instead of attending to all hidden states, local attention attends to a subset of hidden states, thus reducing the computational cost [154]. Google Neural Machine Translation is a popular encoder-decoder model with an attention mechanism that significantly improves the accuracy of machine translation [155]. As shown in Figure 25, the model consists of a multilayer of LSTMs with eight encoder and decoder layers and an attention connection between the bottom layer of the decoder to the top layer of the encoder. Furthermore, to deal with the challenging words to predict, a word is tokenized into subwords e.g. feud is broken down into fe and ud, allowing the model to generalize well to new and uncommon words. A year later, the self-attention mechanism was proposed, significantly improving the overall accuracy of machine translation [16]. Self-attention, also known as intra-attention, allows the deep learning model to capture the dependencies between the input words. The self-attention mechanism is the fundamental building block of the transformer model, which has since become a cornerstone in natural language processing and other domains.
@@
Despite the success of transformer, the model falls short in capturing nuances of human language and struggles with tasks requiring deeper understanding of context. This can be especially challenging when the tasks involve formality, colloquialism, and subtle cultural references that may not directly equivalent in the target language, resulting in inaccurate translation or losing the original meaning. One of the approaches to include context into the input sequence is concatenating the current source sentence with the previous (context) sentences and feeding the whole input to the transformer [156]. The model is trained to predict the translated sentence including the context translation. At inference time, only the translation is considered while the context translation is discarded. Furthermore, the approach encodes the sentence position and segment-shifted position to improve the distinction between current sentences and context sentences. In [157], the source sentence is prefixed with the summary of the document to contextualize the input sentence. The summary is the set of salient words that represents the essence of the document, resolving ambiguity associated with the translation. A study was conducted to determine the optimal technique of aggregating contextual features [158]. Three techniques were studied namely concatenation mode, flat mode and hierarchical mode, and the experimental results show that concatenation mode achieved the best results. In [159], a training method is introduced to train the deep learning machine translation model to generate translation involving honorific words. The training method indicates the honorific context in the target sentence using an honorific classifier to guide the model to attend to the related tokens. Unlike other studies where the context features are included by concatenation, the training method assigns weights to the context tokens indicated by the honorific classifier. This allows the model to generate a more accurate translation with honorifics. Finally, the performance of transformer relies on large-scale training data. However, for the vast majority of languages, only limited amounts of training data exist. To mitigate this problem, recent studies introduce shallow transformer architectures [160], explore the effect of hyperparameter finetuning [161], exploiting monolingual corpus to enhance the bilingual dataset for model training [162] and leveraging visual input as contextual information for the translation task [163].
@@ 4.2.3 Text Generation
Text generation refers to the process of creating texts based on a given input whereby the input can be in the form of texts, images, graphs, tables or even tabular data. Due to the various forms of inputs, text generation has a wide range of applications, including creative writing, image captioning and music generation. This section focuses on the progress made in text-to-text generation tasks such as question answering, dialogue generation and text summarization. The recurrent neural network and its variants play an important role in text generation tasks for their strong ability to model sequential data. One of the earliest works on question answering is based on the RNN-based encoder-decoder model whereby the encoder takes the question embedding and processes it using bi-directional LSTM, and the decoder generates the corresponding answer [164]. Additionally, to prevent semantic loss and enable the model to focus on the important words in the input sequence, a convolution operation is applied to the word embedding, and an attention mechanism is then used to attend to the output of the convolution operation. Similar work is reported in [165] in which a knowledge-based module is introduced to calculate the relevance score between the question and the relevant facts in the knowledge base. This improves the text (answer) generation by the decoder. Another work is described in [166] where an encoder-decoder with attention for dialogue generation is optimized using reinforcement learning. The model is first trained in supervised learning manner and then improved using the policy gradient method to diversify the responses. Ambiguous content in question answering sentences is a challenge in text generation and can lead to incorrect and uncertain responses. Cross-sentence context aware bi-directional model introduces a parallel attention module to compute the co-attention weights at the sentence level, accounting for the relationships and similarities in the question and the answer [167].
@@
The transformer has been leveraged for text generation tasks. An incremental transformer-based encoder is proposed to incrementally encode the historical sequence of conversations [168]. The decoder is a two-pass decoder that is based on the deliberation network, generates the next sentence. The first pass focuses on contextual coherence of the conversations while the second pass refines the output of the first pass. BERT and ALBERT have been used as pre-trained models for question answering task [169]. The study found that the performance of the models is sensitive to random assignment of the initial weights especially on small datasets [170]. T-BERTSum is a model based on BERT, designed to address the challenge of long text dependence and leveraging latent topic mapping in text summarization [171]. The model integrates a neural topic module to infer topics and guide summarization, uses a transformer network to capture long-range dependencies and incorporates multilayers of LSTM for information filtering. Exploiting domain knowledge is essential in reducing the semantic gap between the deep learning models and the text corpus. KeBioSum is a knowledge infusion framework to inject domain knowledge into the pre-trained BERTs for text summarization [172]. In the framework, the relevant information is detected and extracted from the domain knowledge, generating label sequences of the sentences. The label data is then used to train the text summarization model using discriminative and generative training approaches, infusing the knowledge into the model.
@@ 4.3 Time Series and Pervasive Computing
Pervasive computing, often referred to as ubiquitous computing, is the process of integrating computer technology into everyday objects and surroundings so that they become intelligent, networked, and able to communicate with one another to offer improved services and functionalities [173]. According to He et al [174], the role of pervasive computing is foremost in the field where it provides the ability to distribute computational services to the surroundings where people work, leading to trust, privacy, and identity. Examples of pervasive computing applications include smart homes with connected appliances, wearable devices that monitor health and fitness, smart cities with sensor networks for traffic management, and industrial applications that utilize the Internet of Things (IoT) for monitoring and control. Generally, the continuous interaction of interconnected devices in pervasive computing often result in time series data, which captures the evolution of various parameters over time.
For instance, medical sensors, such as electrocardiograms (ECG) and electroencephalograms (EEG), generate time series data that contain critical diagnostic information, which deep learning can use to detect anomalies, predict diseases, and classify medical conditions with improved accuracy. Also, devices such as accelerometers, magnetometers and gyroscopes, among others can be used to capture human activity signals, which are often represented as time series of state changes [175]. In traditional machine learning, features such as mean, variance and others are manually extracted from times series of state changes before human activity classification. However, deep learning models automatically extract features [61]. Also, in other fields such as finance, which entail time series data, deep learning has been instrumental in stock price prediction [176], fraud detection [177], and algorithmic trading [178], among others. Generally, deep learning networks excel at capturing intricate temporal relationships within time-series data, enabling more precise predictions and improved decision-making. Based on this, several deep learning models have been employed for feature learning across various time series and pervasive computing domains.
@@ 4.3.1 Human Activity Recognition
Several researchers have proposed the use of CNN, RNN, and Hybrid models for deep learning based feature learning in wearable sensor HAR. For instance, using two-dimensional CNN (Conv2D), several researchers, as seen in [181], [182] and [183], among others, have developed deep learning models for wearable sensor HAR, despite the time series nature of the data. This is often done by treating the time series signals from wearable sensors as 2D images by reshaping them appropriately. To achieve this, researchers often organize each time series signal into a matrix format, with time along one axis and sensor dimensions along the other, before creating a pseudo-image representation, which allows the matrix to be be fed into Conv2D layers for feature extraction. Conv2D layers excel at capturing spatial patterns and relationships within images, and by treating the time series data as images, these layers can learn relevant spatial features that contribute to activity recognition. The convolution operation performed by Conv2D filters across both the time and sensor dimensions, allowing the network to identify patterns and features that may be indicative of specific activities.
@@
For instance, Ragab et al. [184] proposed a random search Conv1D model, and evaluated the performance of the model on UCI-HAR dataset. The result showed that the model achieved a recognition accuracy of 95.40% when classifying the six activities in the dataset. However, the model exhibited extended training times due to the dynamic nature of some activities within the dataset. To address this, [185] proposed the use of varying kernel sizes in Conv1D layers to recognize various activities, including sitting, standing, walking, sleeping, reading, and tilting. Also, a few Conv1D layers were stacked in order to streamline the time optimization process for training the neural network. Also, some researchers have proposed models that combine machine learning algorithms with Conv1D in HAR, as seen in Shuvo et al. [186]. Their work presented a two-stage learning process to improve HAR by classifying activities into static and dynamic using Random Forest, before using Support Vector Machine to identify each static activity, and Conv1D to recognize dynamic activities. The result showed that the method achieved an accuracy of 97.71% on the UCI-HAR dataset.
@@
Following these advancements, several researchers have further explored Conv1D architectures with various modifications, to enhance feature learning in activity recognition systems. For example, Han et al. [187] developed a two-stream CNN architecture as a plug-and-play module to encode contextual information of sensor time series from different receptive field sizes. The module was integrated into existing deep models for HAR at no extra computation cost. Experiments on OPPORTUNITY, PAMAP2, UCI-HAR and USC-HAD datasets showed that the module improved feature learning capabilities. A similar research reported in [188] proposed the WSense module to address the issue of differences in the quality of features learnt, regardless of the size of the sliding window segmentation, and experimented on PAMAP2 and WISDM datasets. The results showed that by plugging the WSense module into Conv1D architectures, improved activity features can be learned from wearable sensor data for human activity recognition.
@@
Similarly, some researchers have also proposed the use of standalone RNNs in HAR, and a hybrid of Conv1D architectures with RNNs such as LSTMs [189], BiLSTMs [190, 191], GRUs [192] and BiGRUs [193, 194] in order to fully harness the feature learning capabilities of both CNN and RNNs. For instance, Nafea et al. [195], leveraged Bi-LSTM and Conv1D with increasing kernel sizes to learn features at various resolutions. Human activity features were extracted using the stacked convolutional layers with a Bi-LSTM layer, before including a flattening layer and a fully connected layer for subsequent classification. However, the model had issues extracting quality features of dynamic activities compared to static activities. To address such issues, some research works have incorporated attention mechanisms in Conv1D-based architectures to improve feature learning of dynamic and complex activities from time series signals obtained from wearable sensors. For example, Khan and Ahmad [196] designed three lightweight convolutional heads, with each specialized in feature extraction from wearable sensor data. Each head comprised stacked layers of Conv1Ds, along with embedded attention mechanisms to augment feature learning. The results demonstrated that integrating multiple 1D-CNN heads with attention mechanisms can enhance feature learning for Human Activity Recognition. Ayo and Noor [197] designed three feature learning pipelines, each pipeline consisting of two concurrent layers of Conv1D and LSTM with max-pooling, which are then concatenated and processed using an attention mechanism to enhance feature learning. These diverse modifications and adaptations showcase the versatility and potential of deep learning models in achieving state-of-the-art in HAR systems.

### v6

@@ 4.2.2 Neural Machine Translation
Neural machine translation (NMT) refers to the automated process of translating text from one language to another language. Numerous deep learning models have been proposed for NMT, which can be categorized into RNN-based and CNN-based models. One of the first successful RNN-based models is the encoder-decoder [49, 172]. The model consists of two connected subnetworks (the encoder and the decoder) for modelling the translation process, as shown in Figure 28. The encoder reads the source sentence word by word and produces a fixed-length context vector (final hidden state). This process is known as source sentence encoding, as shown in the figure. Given the context vector, the decoder generates the target sentence (translation) word by word. This modelling of the translation can be seen as a mapping between the source sentence to the target sentence via the intermediate context vector in the semantic space. The context vector represents the summary of the input sequence’s semantic meaning, providing a compressed representation that captures the essence of the source sentence. However, the compression process can sometimes result in the loss of information, especially those early in the sequence. Bidirectional RNN may mitigate the loss of information by modelling the sequence in reverse order. However, the problem can persist, particularly in cases where the input is a long sequence.
Figure 28: The architecture of an encoder-decoder. Adapted from [173].
@@
Attention mechanism was introduced to solve the problem of learning long input sequences [174]. Attention alleviates this issue by attending on different words of the input sequences when predicting the target sequences at each time step. Unlike the standard encoder-decoder model, attention derives the context vector from the hidden states of both the encoder and decoder, and the alignment between the source and target. This mechanism allows the model to focus on the important words, increasing the overall accuracy of the translation. Several alignment score functions have been proposed for calculating the attention weights. Some popular functions are additive [174], dot-product, location-based [175], and scaled dot-product [18]. The attention weights are calculated by attending to the entire hidden states of the encoder. This attention, also known as global attention, is computationally expensive. Instead of attending to all hidden states, local attention attends to a subset of hidden states, thus reducing the computational cost [175]. Google Neural Machine Translation is a popular encoder-decoder model with an attention mechanism that significantly improves the accuracy of machine translation [176]. As shown in Figure 29, the model consists of a multilayer of LSTMs with eight encoder and decoder layers and an attention connection between the bottom layer of the decoder to the top layer of the encoder. Furthermore, to deal with the challenging words to predict, a word is tokenized into subwords e.g. feud is broken down into “fe” and “ud”, allowing the model to generalize well to new and uncommon words. A year later, the self-attention mechanism was proposed, significantly improving the overall accuracy of machine translation [18]. Self-attention, also known as intra-attention, allows the deep learning model to capture the dependencies between the input words. The self-attention mechanism is the fundamental building block of the transformer model, which has since become a cornerstone in natural language processing and other domains.
@@
Despite the success of transformer, the model falls short in capturing nuances of human language and struggles with tasks requiring deeper understanding of context. This can be especially challenging when the tasks involve formality, colloquialism, and subtle cultural references that may not directly equivalent in the target language, resulting in inaccurate translation or losing the original meaning. One of the approaches to include context into the input sequence is concatenating the current source sentence with the previous (context) sentences and feeding the whole input to the transformer [177]. The model is trained to predict the translated sentence including the context translation. At inference time, only the translation is considered while the context translation is discarded. Furthermore, the approach encodes the sentence position and segment-shifted position to improve the distinction between current sentences and context sentences. In [178], the source sentence is prefixed with the summary of the document to contextualize the input sentence. The summary is the set of salient words that represents the essence of the document, resolving ambiguity associated with the translation. A study was conducted to determine the optimal technique of aggregating contextual features [179]. Three techniques were studied namely concatenation mode, flat mode and hierarchical mode, and the experimental results indicate that concatenation mode achieved the best results. In [180], a training method is introduced to train the deep learning machine translation model to generate translation involving honorific words. The training method indicates the honorific context in the target sentence using an honorific classifier to guide the model to attend to the related tokens. Unlike other studies where the context features are included by concatenation, the training method assigns weights to the context tokens indicated by the honorific classifier. This allows the model to generate a more accurate translation with honorifics.
The performance of transformers relies on large-scale training data. However, for the vast majority of languages, only limited amounts of training data exist. To mitigate this problem, recent studies introduce shallow transformer architectures [181], explore the effect of hyperparameter finetuning [182] and leveraging visual input as contextual information for the translation task [183]. The semi-supervised neural machine translation is used for translating low resource Egyptian Arabic dialects to modern standard Arabic [184]. The study utilizes three datasets: bilingual Egyptian-standard Arabic and two monolingual Egyptian Arabic and standard Arabic. First, a transformer-based model is trained in a supervised manner using the bilingual dataset. Then, the model is trained in an unsupervised manner using both monolingual datasets. The unsupervised approach employs an Encoder-Decoder model with Byte Pair Encoding for tokenization and handling unknown words. The monolingual corpora are merged and used to improve the model by iteratively generating synthetic sentence pairs between Egyptian Arabic and standard Arabic, allowing the model to learn the correspondence between them. In [185], the authors exploit monolingual corpus to enhance the bilingual dataset for model training. Furthermore, a new loss function is proposed as a replacement for traditional cross-entropy loss, allowing the model to learn with uncertainty in the presence of noise. Additionally, contrastive re-ranking is employed to refine translation results by selecting the most confident output from multiple candidates. Table 7 lists the summary of state-of-the-art neural machine translation.
Table 7: Summary of state-of-the-art neural machine translation.
Ref. Description Results (Datasets and Metrics)
[…]
ContraPro: 82.54
[…]
Voita: 75.94
[…]
BLEU: 22.0, COMET: 0.785
[…]
BLEU: 30.89
[…]
BLEU: 36.84
[…]
BLEU: 20.06
[…]
BLEU: 20.50
[…]
BLEU: 27.90
[…]
BLEU: 20.70
[…]
BLEU: 13.67
[184], 2024 This study employed a semi-supervised neural machine translation approach to translate low-resource Egyptian Arabic dialects to Modern Standard Arabic, utilizing bilingual and monolingual datasets, with a transformer-based model trained in both supervised and unsupervised manners to generate synthetic sentence pairs and improve translation accuracy Egyptian-standard Arabic
BLEU: 24 (300 word embedding)
BLEU: 29.5 (512 word embedding)
[…]
BLEU: 28.12
[…]
BLEU: 23.53
[…]
BLEU: 28.91
[…]
BLEU: 22.76
@@ 4.2.3 Text Generation
Text generation refers to the process of creating texts based on a given input, whereby the input can be in the form of texts, images, graphs, tables or even tabular data. Due to the various forms of inputs, text generation has a wide range of applications, including creative writing, image captioning and music generation. This section focuses on the progress made in text-to-text generation tasks such as question answering, dialogue generation and text summarization. The recurrent neural network and its variants play an important role in text generation tasks for their strong ability to model sequential data. One of the earliest works on question answering is based on the RNN-based encoder-decoder model, whereby the encoder takes the question embedding and processes it using bidirectional LSTM, and the decoder generates the corresponding answer [186]. Additionally, to prevent semantic loss and enable the model to focus on the important words in the input sequence, a convolution operation is applied to the word embedding, and an attention mechanism is then used to attend to the output of the convolution operation. Similar work is reported in [187] in which a knowledge-based module is introduced to calculate the relevance score between the question and the relevant facts in the knowledge base. This improves the text (answer) generation by the decoder. Another work is described in [188] where an encoder-decoder with attention for dialogue generation is optimized using reinforcement learning. The model is first trained in a supervised learning manner and then improved using the policy gradient method to diversify the responses. Ambiguous content in question answering sentences is a challenge in text generation and can lead to incorrect and uncertain responses. Cross-sentence context aware bidirectional model introduces a parallel attention module to compute the co-attention weights at the sentence level, accounting for the relationships and similarities in the question and the answer [189].
@@
The transformer has been leveraged for text generation tasks. An incremental transformer-based encoder is proposed to incrementally encode the historical sequence of conversations [190]. The decoder is a two-pass decoder that is based on the deliberation network, generates the next sentence. The first pass focuses on contextual coherence of the conversations, while the second pass refines the output of the first pass. BERT and ALBERT have been used as pre-trained models for question answering task [191]. The study found that the performance of the models is sensitive to random assignment of the initial weights especially on small datasets [192]. T-BERTSum is a model based on BERT, designed to address the challenge of long text dependence and leveraging latent topic mapping in text summarization [193]. The model integrates a neural topic module to infer topics and guide summarization, uses a transformer network to capture long-range dependencies and incorporates a multilayer of LSTM for information filtering.
The generated texts often lack diversity and may exhibit repetitive patterns. To address this issue, feature-aware conditional GAN (FA-GAN) is proposed for controllable category text generation [194]. The generator consists of BERT, a category encoder, a relational memory core (RMC) decoder. BERT acts as a feature encoder, improves contextual representation and mitigates mode collapse while the category encoder embeds categorical information for text generation. The RMC decoder utilizes a self-attention mechanism to capture interactions between features, generating more expressive and diverse texts. The discriminator includes an additional classification head to ensure the generated texts match specified categories. Traditional text generation has largely focused on binary style transfers, but real-world applications require capturing diverse styles, which existing methods struggle to achieve. In [195], a multi-class conditioned text generation model is proposed using a transformer-based decoder with an adversarial module, a style attention module and a generation module. The adversarial module extracts style-excluded representation from the input text. The style attention mechanism then introduces the desired style into the text representation through concat attention mechanism, producing a conditioned representation. Finally, the generation module utilizes the conditioned representation to generate stylistically diverse text.
Exploiting domain knowledge is essential in reducing the semantic gap between the deep learning models and the text corpus. KeBioSum is a knowledge infusion framework to inject domain knowledge into the pre-trained BERTs for text summarization [196]. In the framework, the relevant information is detected and extracted from the domain knowledge, generating label sequences of the sentences. The label data is then used to train the text summarization model using discriminative and generative training approaches, infusing the knowledge into the model. Large language models (LLMs) have been used for text generation with exceptional quality and diversity. LLMs, trained on extensive corpus data, have a deep understanding of human language, allowing them to interpret and generate texts. In [197], GPTGAN is introduced that leverages an LLM as a guiding mentor to a GAN-Autoencoder model for text generation. The approach involves the GPT model to generate a sequence of words given a subset of the input sequence, and the generated text is mapped into latent space by the transformer-based encoder. The latent representation is then used by the transformer-based decoder to generate the text. Furthermore, local discriminators are introduced to refine the text generation because the generated words by the GPT model may be inaccurate. Table 8 lists the summary of state-of-the-art text generation.
Table 8: Summary of state-of-the-art text generation.
Ref. Description Results (Datasets and Metrics)
[193], 2021 The paper proposed a text summarization model that integrates neural topic model with BERT to guide the text summarization based on inferred topics. It leverages transformers for long-term dependencies and combines LSTM layers with a gated network to enhance extractive and abstractive summarization of social texts CNN/Daily Mail
Rouge-1: 43.58, Rouge-2: 20.45, Rouge-L: 34.60
Xsum
Rouge-1: 39.90, Rouge-2: 17.48, Rouge-L: 29.85
[194], 2023 This study proposed FA-GAN featuring BERT for contextual representation, a category encoder for embedding categorical information, and an RMC decoder for generating diverse texts, while the discriminator ensures category consistency MR-10
BLEU-2: 0.560
MR-20
BLEU-2: 0.674
AM-30
BLEU-2: 0.748
[…]
self-BLEU: 0.257, S-ACC: 0.375
[…]
self-BLEU: 0.096, S-ACC: 0.191
[…]
self-BLEU: 0.257, S-ACC: 0.198
[…]
self-BLEU: 0.081, S-ACC: 0.207
*S-ACC (style accuracy)
[196], 2022 The paper proposes a knowledge-based summarization model that integrates medical knowledge into the pretrained models using a lightweight knowledge adapter. It employs generative and discriminative training to predict and reconstruct PICO elements, enhancing domain-specific text summarization CORD-19
Rouge-1: 32.04, Rouge-2: 12.61, Rouge-L: 29.10
PubMed-Long
Rouge-1: 36.39, Rouge-2: 16.27, Rouge-L: 33.28
[197], 2025 This study proposed GPTGAN, an approach that enhances adversarial text generation using a GPT model as a mentor to the GAN model and employing a multiscale discriminator framework to balance text quality and diversity MSCOCO
BLEU: 98.7
WMTNews
BLEU: 93.4
Persian COCO
BLEU: 97.9
@@ 4.3 Time Series and Pervasive Computing
Pervasive computing, often referred to as ubiquitous computing, is the process of integrating computer technology into everyday objects and surroundings so that they become intelligent, networked, and able to communicate with one another to offer improved services and functionalities [198]. According to He et al. [199], the role of pervasive computing is foremost in the field where it provides the ability to distribute computational services to the surroundings where people work, leading to trust, privacy, and identity. Examples of pervasive computing applications include smart homes with connected appliances, wearable devices that monitor health and fitness, smart cities with sensor networks for traffic management, and industrial applications that utilize the Internet of Things (IoT) for monitoring and control. Generally, the continuous interaction of interconnected devices in pervasive computing often results in time series data, which captures the evolution of various parameters over time.
For instance, medical sensors, such as electrocardiograms (ECG) and electroencephalograms (EEG), generate time series data that contain critical diagnostic information, which deep learning can use to detect anomalies, predict diseases, and classify medical conditions with improved accuracy. Furthermore, devices such as accelerometers, magnetometers and gyroscopes, among others, can be used to capture human activity signals, which are often represented as time series of state changes [200]. In traditional machine learning, features such as mean, variance, and others are manually extracted from times series of state changes before human activity classification. However, deep learning models automatically extract features [78]. Furthermore, in other fields such as finance, which entail time series data, deep learning has been instrumental in stock price prediction [201], fraud detection [202], and algorithmic trading [203], among others. Generally, deep learning networks excel at capturing intricate temporal relationships within time-series data, enabling more precise predictions and improved decision-making. Based on this, several deep learning models have been employed for feature learning across various time series and pervasive computing domains.
@@ 4.3.1 Human Activity Recognition
Several researchers have proposed the use of CNN, RNN, and Hybrid models for deep learning-based feature learning in wearable sensor HAR. For instance, using two-dimensional CNN (Conv2D), several researchers, as seen in [206], [207] and [208], among others, have developed deep learning models for wearable sensor HAR, despite the time series nature of the data. This is often done by treating the time series signals from wearable sensors as 2D images by reshaping them appropriately. To achieve this, researchers often organize each time series signal into a matrix format, with time along one axis and sensor dimensions along the other, before creating a pseudo-image representation, which allows the matrix to be fed into Conv2D layers for feature extraction. Conv2D layers excel at capturing spatial patterns and relationships within images, and by treating the time series data as images, these layers can learn relevant spatial features that contribute to activity recognition. The convolution operation performed by Conv2D filters across both the time and sensor dimensions, allowing the network to identify patterns and features that may be indicative of specific activities.
@@
For instance, Ragab et al. [209] proposed a random search Conv1D model, and evaluated the performance of the model on UCI-HAR dataset. The result showed that the model achieved a recognition accuracy of 95.40% when classifying the six activities in the dataset. However, the model exhibited extended training times due to the dynamic nature of some activities within the dataset. To address this, [210] proposed the use of varying kernel sizes in Conv1D layers to recognize various activities, including sitting, standing, walking, sleeping, reading, and tilting. Furthermore, a few Conv1D layers were stacked to streamline the time optimization process for training the neural network. In [211], a signal segmentation method is proposed based on similarity between two subsequent windows. The method can not only achieve more accurate segmentation, but also distinguish between transitional and non-transitional windows. Based on the distinction, two deep learning models with convolutional and fully connected layers are trained: one for transitional activities and the other for non-transitional activities. In [212], signal segmentation is treated as a binary classification task, distinguishing windows as either transitional or non-transitional. To this end, a deep learning model with parallel Conv1D pipelines is proposed to capture temporal dependencies within the window sequence. Furthermore, some researchers have proposed models that combine machine learning algorithms with Conv1D in HAR, as seen in Shuvo et al. [213]. Their work presented a two-stage learning process to improve HAR by classifying activities into static and dynamic using Random Forest, before using Support Vector Machine to identify each static activity, and Conv1D to recognize dynamic activities. The result shows that the method achieved an accuracy of 97.71% on the UCI-HAR dataset.
@@
Following these advancements, several researchers have further explored Conv1D architectures with various modifications, to enhance feature learning in activity recognition systems. For example, Han et al. [214] developed a two-stream CNN architecture as a plug-and-play module to encode contextual information of sensor time series from different receptive field sizes. The module was integrated into existing deep models for HAR at no extra computation cost. Experiments on OPPORTUNITY, PAMAP2, UCI-HAR and USC-HAD datasets show that the module improved feature learning capabilities. A similar research is reported in [215] proposed the WSense module to address the issue of differences in the quality of features learnt, regardless of the size of the sliding window segmentation, and experimented on PAMAP2 and WISDM datasets. The results show that by plugging the WSense module into Conv1D architectures, improved activity features can be learned from wearable sensor data for human activity recognition.
@@
Hybrid Models: Researchers have also proposed the use of standalone RNNs in HAR, and a hybrid of Conv1D architectures with RNNs such as LSTMs [216], BiLSTMs [217, 218], GRUs [219] and BiGRUs [220] to fully harness the feature learning capabilities of both CNN and RNNs. For instance, Nafea et al. [221], leveraged Bi-LSTM and Conv1D with increasing kernel sizes to learn features at various resolutions. Human activity features were extracted using the stacked convolutional layers with a Bi-LSTM layer, before including a flattening layer and a fully connected layer for subsequent classification. However, the model had issues extracting quality features of dynamic activities compared to static activities. To address such issues, some research works have incorporated attention mechanisms in Conv1D-based architectures to improve feature learning of dynamic and complex activities from time series signals obtained from wearable sensors. For example, Khan and Ahmad [222] designed three lightweight convolutional heads, with each specialized in feature extraction from wearable sensor data. Each head comprised stacked layers of Conv1Ds, along with embedded attention mechanisms to augment feature learning as shown in Figure 30. The results demonstrated that integrating multiple 1D-CNN heads with SE attention can enhance feature learning for Human Activity Recognition.
Ayo and Noor [223] designed three feature learning pipelines, each pipeline consisting of two concurrent layers of Conv1D and LSTM with maximum pooling, which are then concatenated and processed using a channel-wise attention mechanism to enhance feature learning. In [224], a sequential channel-temporal attention is proposed for multi-modal activity recognition. The channel attention is similar to SE attention, but with average pooling and maximum pooling pipelines to squeeze to temporal dimension. The pooled features are then combined through addition operation. As for the temporal attention, average pooling and maximum pooling are applied along the channel dimension, producing two pooled features which are then combined through concatenation. The attention module is incorporated into the residual blocks of ResNet-like model to improve feature extraction. Similar work is reported in [225] whereby a hybrid sequential channel-spatial attention is proposed for a lightweight activity recognition model. The lightweight model consists of four convolutional blocks integrated with the attention module and two LSTM layers. Although the proposed attentions have shown a strong potential in improving model performance, their sequential processing inherently prioritize one of the feature map dimensions, impeding the model’s ability to capture holistic feature representations. In [226], a triple cross-domain attention is proposed to blend three attention branches to improve feature extraction in HAR. For each attention branch, z-pooling [227] and convolution operations are applied along the corresponding dimension to generate the attention weights. Then, the feature maps are combined via average operation. The model consists of three residual blocks where each block is integrated with the attention module. These diverse modifications and adaptations showcase the versatility and potential of deep learning models in achieving state-of-the-art in HAR systems.